OpenStudy (anonymous):

Kareem owes $5,127 on a credit card with a 15.9% interest rate compounded monthly. What is the monthly payment he should make in order to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? $79.94 $427.25 $464.93 $982.42

terenzreignz (terenzreignz):

Annuities? :D

terenzreignz (terenzreignz):

That's one nasty interest-rate... either unrealistic or just plain evil :P @LuluScavino Are you still here?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

So, the annuity involves a fixed monthly payment, let's call that x, paid over 12 months, with its rate of interest being 15.9% compounded monthly... And we want its present value to be 5127...

OpenStudy (anonymous):

okay so what do i do

terenzreignz (terenzreignz):

Well, if the interest rate is 15.9% *compounded monthly* How much is the effective rate of interest per month?

OpenStudy (anonymous):

im so confused sorry?

terenzreignz (terenzreignz):

15.9% annual interest *compounded monthly* That means... how much is the interest per month?

OpenStudy (anonymous):

1.325

terenzreignz (terenzreignz):

That's right. Don't forget the % sign. So, the annuity formula goes... \[\Large a_{\bar{n|}}=\frac{1-(1+r)^{-n}}{r}\] Where r is the effective rate of interest per period...

terenzreignz (terenzreignz):

Now the present value is given by... \[\Large x\cdot a_{\bar{n|}}\] And this should be equal to 5127. Can you now solve for x?

OpenStudy (anonymous):

5127(a)

terenzreignz (terenzreignz):

No... 5127 is the present value and 5127 = ax Anyway, you'd do well to solve the annuity first.. what is a? Remember, your r is 1.325%

OpenStudy (anonymous):

oh .....

terenzreignz (terenzreignz):

a is not your answer okay? But it will lead you to it almost instantly. Now, solve for a...

terenzreignz (terenzreignz):

\[\Large a_{\bar{n|}}=\frac{1-(1+r)^{-n}}{r}\] But specifically, solve \[\Large a_{\bar{12|}}\]

OpenStudy (anonymous):

but im confused on the formula how does it all come together?

terenzreignz (terenzreignz):

Patience. First, I want you to solve for \(\Large a_{12}\)...

OpenStudy (anonymous):

im so sorry i honestly need to do this quicker because i have lots of questions to answer and this is confusing more than before

OpenStudy (anonymous):

i really appreciate you for your help but can you help me quicker please please

terenzreignz (terenzreignz):

Well, it is rather quick, just key it into the calculator... \[\Large a_{\bar{12|}}=\frac{1-(1+r)^{-12}}{r}\] we have n = 12, because 12 months will pass... now, just key it in the calculator, and tell me what you get for \(\Large a_{12}\)

OpenStudy (anonymous):

i dont where to find that on my calc

OpenStudy (anonymous):

im so so sorry u must hate me

terenzreignz (terenzreignz):

Where to find what?

terenzreignz (terenzreignz):

It's just addition, subtraction, division, an exponent... what can't you find? lol

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

I'm sorry, what?

OpenStudy (anonymous):

Idk it

terenzreignz (terenzreignz):

\[\Large a_{\bar{12|}}=\frac{1-(1+\color{red}r)^{-12}}{\color{red}r}\] What is r?

OpenStudy (anonymous):

1.325%

terenzreignz (terenzreignz):

or, in decimal form?

OpenStudy (anonymous):

0.1325

terenzreignz (terenzreignz):

no... 0.01325 So now you have the value of r... now I trust you'll no longer have any difficulty solving \[\Large a_{\bar{12|}}=\frac{1-(1+r)^{-12}}{r}\]?

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

63.4

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

nope :)

OpenStudy (anonymous):

so then?

OpenStudy (anonymous):

is the answer c?

terenzreignz (terenzreignz):

I don't know, let's find out :)

terenzreignz (terenzreignz):

And to find out, it is absolutely crucial for you to find \(a_{12}\)

OpenStudy (anonymous):

how?

terenzreignz (terenzreignz):

By following this formula...

terenzreignz (terenzreignz):

\[\Large a_{\bar{12|}}=\frac{1-(1+r)^{-12}}{r}\]

terenzreignz (terenzreignz):

Here, I'll even replace r immediately... \[\Large a_{\bar{12|}}=\frac{1-(1+0.01325)^{-12}}{0.01325}\]

terenzreignz (terenzreignz):

It really is just a matter of keying it into the calculator (correctly!)

OpenStudy (anonymous):

2.5

terenzreignz (terenzreignz):

Not even close, @LuluScavino Are you guessing? lol... naughty naughty...

OpenStudy (anonymous):

no thats what i got

terenzreignz (terenzreignz):

Well, it's not right, so you must have made an error in typing into the calculator...

OpenStudy (anonymous):

okay so then what would it be please please please help me get this

terenzreignz (terenzreignz):

I'm trying :) But you have to help me help you ^_^ You'd be in a rather sorry shape if you can't key this \[\Large a_{\bar{12|}}=\frac{1-(1+0.01325)^{-12}}{0.01325}\] correctly into your calculator ... So please, try again, and be extra careful...

OpenStudy (anonymous):

140

terenzreignz (terenzreignz):

The problem with this calculator shenanigan is that I can't see where you made your error... and on top of that, we might not even be using the same kind of calculator :) and no, 140 is not correct...

OpenStudy (anonymous):

then can you please just say it

terenzreignz (terenzreignz):

That's like saying the answer...an act heavily discouraged :) Besides, I have faith in you ^_^

OpenStudy (anonymous):

but ive tried everything and idk what to do?

OpenStudy (anonymous):

please i beg u are the last question i have

terenzreignz (terenzreignz):

No you haven't... because... you haven't tried doing it correctly yet :) Don't give up so easily...

terenzreignz (terenzreignz):

What calculator are you using, btw?

OpenStudy (anonymous):

casio

terenzreignz (terenzreignz):

model...?

OpenStudy (anonymous):

fx 300 ms

OpenStudy (anonymous):

two way power

terenzreignz (terenzreignz):

Then I suggest you do it one at a time... work out the numerator first... \[\Large a_{\bar{12|}}=\frac{\color{blue}{1-(1+0.01325)^{-12}}}{0.01325}\] And when it's done, just divide everything by 0.01325

OpenStudy (anonymous):

please tell me its -63.444

terenzreignz (terenzreignz):

Impossible.. numerator must be positive...

OpenStudy (anonymous):

ok ive tried enough

OpenStudy (anonymous):

SERIOUSLY

OpenStudy (anonymous):

11.02

terenzreignz (terenzreignz):

ohh... finally :)

terenzreignz (terenzreignz):

Okay so we have \(\Large a_{12}=11.02\) And \[\Large x \cdot a_{12}=5127\] Now it's just a simple matter of solving for x

OpenStudy (anonymous):

okay hold up

OpenStudy (anonymous):

i was rights its C

terenzreignz (terenzreignz):

Yeah... but I suspected you were guessing... so I put you up to it first :)

OpenStudy (anonymous):

well look who proved u wrong

OpenStudy (anonymous):

Thanks for all the torture :) lol

terenzreignz (terenzreignz):

You didn't... If you weren't guessing, you wouldn't have had any difficulty whatsoever with \(\Large a_{12}\) Prove me wrong next time :P ... If I'm ever wrong >:D

OpenStudy (anonymous):

Lol (: so then i can ask more questions tmw

terenzreignz (terenzreignz):

We shall see... It's the weekend after all, and let's face it, who studies during weekends?

OpenStudy (anonymous):

me

terenzreignz (terenzreignz):

Then you're not spending your weekends correctly :P

OpenStudy (anonymous):

adios